# Pixel-Aligned Expressive Human Mesh Recovery

### v1

@@ 6 Conclusion
Appendix A Our Motivation
@@
During this process, we identify two key challenges of this strategy. First, existing data preprocessing pipelines jointly handle face, hand, and body parameters, leading to limited annotation accuracy and reduced data diversity (e.g., a lack of close-up facial images). We therefore redesign the data processing pipeline to provide higher-quality supervision (Sec. 3.4). Second, we observe that a simple ViT struggles to capture fine-grained details from low-resolution inputs, motivating us to introduce pixel-level supervision for further refinement. As a result, we achieve an efficient and accurate model that enables real-time estimation of both SMPLX and FLAME parameters, supporting downstream applications. We are the first to explore a simple ViT-based architecture for extracting full-dimensional human body parameters and to validate the feasibility of this approach, achieving fast inference at 100 FPS and aiming to provide a fast and accurate interface for downstream applications.
@@ Appendix B Training Data Generation: Part-level pseudo-label refinement strategy
For supervision data annotation, we design a novel part-level pseudo-label refinement pipeline, as shown in Fig. 17. Given a single human image, we first extract coarse parameters for facial expressions, body pose, and hand pose using three different models: TEASER, ProHMR, and Hamer [47]. The extracted parameters are then refined under the supervision of 2D keypoints predicted by the DWPOSE [75] model. As a result, we obtain four sets of parameters: FLAME, scale, SMPL, and SMPL-X (hand). For SMPL parameters, we additionally perform a mapping from SMPL to SMPL-X to obtain SMPL-X (body) parameters. This process ultimately provides accurate SMPL-X and FLAME parameters, along with the corresponding scale values.
For learning-based methods, a well-designed annotation pipeline is critical. Our modular annotation procedure is thus an indispensable component for enabling the effective training and performance of our model.
B.1 More Results
@@
Table 7: We primarily compare our method with sota SMPLX–based approaches. Quantitative results on PCK and MPJPE demonstrate that our method achieves more accurate human pose estimation. 1 Multi-HMR always detects additional persons in the PoseTrack dataset; therefore, we do not report its results. 2 test on L40S
Methods Model Predict parts FPS 2 COCO LSP-extend PoseTrack 3DPW
[…]
CLIFF [31] SMPL Body - 0.64 0.88 0.32 0.66 0.75 0.92 - -
PARE [26] SMPL Body - 0.72 0.91 0.27 0.66 0.79 0.93 82.0 50.9
HMR2 [11] SMPL Body - 0.87 0.97 0.53 0.82 0.90 0.98 81.3 54.3
HSMR [67] SKEL Body - 0.87 0.96 0.51 0.81 0.90 0.98 81.5 54.8
SAM3D-body [74] MHR Body, hand 1-2 0.87 - 0.69 - - - 54.8 -
OSX [33] SMPLX Body,face,hand 25 0.70 0.87 0.42 0.73 0.82 0.90 74.7 45.1
Multi-HMR1[2] SMPLX Body,face,hand 10 0.65 0.82 0.43 0.70 - - 78.0 45.9
SMPLest-X [76] SMPLX Body,face,hand 25 0.71 0.91 0.40 0.74 0.82 0.91 74.8 45.8
Ours EHM Body,face,hand 100 0.81 0.94 0.55 0.81 0.87 0.97 71.3 45.3

### v2

@@ 6 Conclusion
Appendix A Method details
A.1 Our Motivation
@@
During this process, we identify two key challenges of this strategy. First, existing data preprocessing pipelines jointly handle face, hand, and body parameters, leading to limited annotation accuracy and reduced data diversity (e.g., a lack of close-up facial images). We therefore redesign the data processing pipeline to provide higher-quality supervision. Second, we observe that a simple ViT struggles to capture fine-grained details from low-resolution inputs, motivating us to introduce pixel-level supervision for further refinement. As a result, we achieve an efficient and accurate model that enables real-time estimation of both SMPLX and FLAME parameters, supporting downstream applications. We are the first to explore a simple ViT-based architecture for extracting full-dimensional human body parameters and to validate the feasibility of this approach, achieving fast inference at 100 FPS and aiming to provide a fast and accurate interface for downstream applications.
A.2 Kinematic Consistency of EHM-s
Here, we describe in detail how the EHM-s model adheres to plausible human body structure. EHM-s is constructed by combining SMPL-X with a scaled FLAME head. The alignment procedure is performed as follows:
1. We first obtain a refined, expression-aware head mesh by passing the predicted shape and expression parameters through FLAME.
2. Next, we generate the SMPL-X template using the corresponding SMPL-X shape and expression parameters.
3. In the canonical space of SMPL-X, we perform joint-based registration to align the head mesh, replace the corresponding head region, and obtain the SMPL-X template with the integrated head.
4. Finally, the updated template from the previous step is processed through LBS to produce the final fused mesh.
The corresponding pseudo-code is presented in Tab. 1
Figure 17: The EHM-s forward pipeline.
Figure 18: Effect of different EHM-s parameters on the resulting mesh.
Algorithm 1 FLAME–SMPL-X Head Fusion Pipeline
Input: Head pose parameters, body pose parameters, FLAME head template, head scale $s$, SMPL-X body template
Output: Fused full-body mesh vertices
Step 1: Compute FLAME head mesh via LBS
$(\mathbf{V}_{h},\mathbf{J}_{h})\leftarrow\mathrm{LBS}(\text{head\_pose},\text{template\_head\_vertices})$ $\triangleright$ Global pose removed
Step 2: Align FLAME head to SMPL-X body
$\mathbf{c}_{h}\leftarrow\mathrm{Mean}(\mathbf{J}_{h}[3{:}5])$
$\mathbf{c}_{b}\leftarrow\mathrm{Mean}(\mathbf{J}_{b}[23{:}25])$
$\hat{\mathbf{V}}_{h}\leftarrow\mathbf{V}_{h}*s-\mathbf{c}_{h}+\mathbf{c}_{b}$
Step 3: Replace SMPL-X head region
$\mathbf{V}_{b}[\text{smplx2flame\_indices}]\leftarrow\hat{\mathbf{V}}_{h}$
Step 4: Final LBS for full-body mesh
$(\mathbf{V},\mathbf{J})\leftarrow\mathrm{LBS}(\text{body\_pose},\text{global\_pose},\mathbf{V}_{b})$
A.3 EHM-s forward process
As shown in Fig. 18, we first start from the SMPL-X basis and apply blend shape deformations using the body shape parameters $\beta_{b}$ and expression parameters $\phi$ to obtain a coarse body template. Next, based on the FLAME basis, we generate accurate and highly expressive facial geometry using the facial shape $\beta_{h}$, expression $\phi_{h}$, and pose $\theta_{h}$ parameters. Since the head root of FLAME is aligned with that of SMPL-X, the head geometry can be replaced and scaled around this shared root to produce a unified and accurate human geometry. Finally, the resulting geometry is posed using the SMPL-X linear blend skinning (LBS) function, yielding the final human mesh.
@@ Appendix B Training Data Generation: Part-level pseudo-label refinement strategy
To construct reliable supervision signals for training, we design a novel part-level pseudo-label refinement pipeline, as illustrated in Fig. 19. Given a single RGB image containing a human subject, we first estimate coarse parametric representations for different body parts using three off-the-shelf models specialized for distinct components: TEASER [35] for facial expression parameters, ProHMR [28] for full-body pose and shape, and HaMeR [46] for detailed hand pose estimation. These models operate independently and provide initial, part-specific predictions that may suffer from inconsistencies and accumulated errors.
To further improve the accuracy and cross-part consistency of these estimates, we refine all extracted parameters under the supervision of 2D keypoints predicted by the DWPOSE [73] model. Specifically, the predicted 2D keypoints serve as geometric constraints to align the projected model joints with image observations, enabling effective correction of pose ambiguities and scale discrepancies.
Through this refinement process, we obtain four sets of optimized parameters, including FLAME parameters for facial geometry and expression, global scale factors, SMPL parameters for the body, and SMPL-X (hand) parameters for articulated hand poses. For body representations, we further convert the refined SMPL parameters into the SMPL-X formulation using a standard parameter mapping, resulting in consistent SMPL-X (body) parameters that are compatible with the hand and face models.
Ultimately, this pipeline produces accurate and coherent SMPL-X and FLAME pseudo-labels, together with reliable scale estimates, which serve as high-quality supervision for subsequent model training. For learning-based methods, a well-designed annotation pipeline is critical. Our modular annotation procedure is thus an indispensable component for enabling the effective training and performance of our model.
@@
B.1 Implementation and training details.
Our training objective consists of a combination of parameter-level supervision and geometric consistency losses. Specifically, we apply a head parameter loss to supervise fine-grained facial representations, including eye pose, jaw articulation, eyelid motion, expression, and shape parameters, using an L1 loss masked by the availability of FLAME annotations. For full-body modeling, we introduce a body parameter loss that supervises SMPL/SMPL-X parameters, including body pose, hand pose, shape, expression, and scale. Body and hand poses are converted from axis-angle representations to rotation matrices before computing an MSE loss, which improves stability in pose regression. The supervision is selectively applied using validity masks to handle incomplete annotations.
In addition to parameter regression, we incorporate 2D keypoint reprojection loss and 3D keypoint loss to enforce geometric consistency. The 2D loss aligns projected joints with image observations using confidence-weighted L1 distances, while the 3D loss supervises root-relative joint positions to preserve plausible human body structures. The final loss is computed as a weighted sum of all components, ensuring balanced optimization across facial details, articulated body motion, and global geometric alignment.
B.2 More Results
We further provide additional qualitative visualizations for analysis, as shown in Fig. 20. The results illustrate that our method is capable of producing reliable reconstructions under diverse and complex poses. In particular, the inferred geometry remains stable and accurate when viewed from side and top perspectives, demonstrating consistent spatial reasoning across different viewpoints.